Express M as a linear combination of the matrices A, B, C, where M = [

4 7

7 9

] , A = [

1 1

1 1

] , B


= [

1 2

3 4

] , C = [

1 1

4 5

] .


1
Expert's answer
2021-11-07T16:41:50-0500

To show M as linear combination s of A B C.. we take three variable a,b,c

And express as follows:


"M=\begin{bmatrix}4\\7\\
7\\
9
\end{bmatrix}=a
\begin{bmatrix}
1\\
1\\
1\\
1
\end{bmatrix}
+b
\begin{bmatrix}
1\\
2\\
3\\
4
\end{bmatrix}
+c
\begin{bmatrix}
1\\
1\\
4\\
5
\end{bmatrix}"



Then we get 4 linear equations


4=1a+1b+1c...........1

7=1a+2b+1c...........2

7=1a+3b+4c............3

9=1a+4b+5c...........4


To solve:

Take eqn. 1 and 2

Subtracting 1 from 2 .


We get b=3.


Now putting value of b in eqn. 2 and 3 ..

Now subtracting 2 from 3.

We get value of c =-1


Now putting values of b and c in equation 4.

We get value of a= 2.

Thus,


By solving these 4 equation we get,

a=2

b=3

c=-1


Now ;

Putting back these values in matrix equation



"M=\begin{bmatrix}4\\7\\
7\\
9
\end{bmatrix}=2
\begin{bmatrix}
1\\
1\\
1\\
1
\end{bmatrix}
+3
\begin{bmatrix}
1\\
2\\
3\\
4
\end{bmatrix}
-1
\begin{bmatrix}
1\\
1\\
4\\
5
\end{bmatrix}"

Thus we get expression for M as linear combination of A,B,C matrices.